OpenStudy (anonymous):

one pound is about 2.2 kilograms. jefferson's dog weights 40 lb.what is the mass of his dog in kilograms?

OpenStudy (anonymous):

1 pound = 2.2 kg 40 * 1 pounds = 40 * 2.2 kg 40 pounds = 88 kg :)

OpenStudy (anonymous):

thank u u a hero :)